OpenStudy (anonymous):

Which expression is equivalent to (1+9x)10

OpenStudy (anonymous):

just expand. 1 times 10 + 9x times 10

OpenStudy (anonymous):

10+90x

OpenStudy (anonymous):

yes, @shirleyx has the correct answer.

OpenStudy (anonymous):

Is 10 in power or 10 in multiplication??

OpenStudy (anonymous):

you need to multiply or it is in power????

OpenStudy (anonymous):

which expression means do you have the options??????

OpenStudy (unklerhaukus):

\[(1+9x)^{10}=\left(1^{10}+\cdots\cdots+(9x)^{10}\right)=\left(1+\cdots\cdots+9^{10}x^{10}\right)\]